# Surrogate Modeling of Trajectory Map-matching in Urban Road Networks using Transformer Sequence-to-Sequence Model

### v2

@@ V-B3 Transformer-based surrogate map-matching
The configuration of our deep-learning surrogate model is set as follows: The embedding dimension ($d_{emb}$) is set to 512, while the MLP intermediate dimension ($d_{ff}$) is set to 1024. We use $N=4$ transformer blocks for both the encoder and decoder, and each multi-head attention block employs $h=4$ linear transformations, resulting in $d_{k}=128$. The Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function is utilized. The batch size is set to 32, and the model trained for a total of 200 epochs. The inference time for map-matching using the constructed transformer model varies with trajectory length, with an average inference time of 0.30s and a standard deviation of 0.16s for trajectories ranging from 6 to 204 road segments.
@@ VI-A2 RNN-based encoder-decoder with attention mechanism
The main components of the RNN-based encoder-decoder architecture are two layers: the embedding layer and the RNN layer. Therefore, the main difference between the RNN-based encoder-decoder with the transformer-based encoder-decoder is its recurrent layer versus the self-attention layer in the transformer. The architecture of the RNN-based encoder-decoder model can be found in [61]. In an RNN-based encoder-decoder, during the encoding, the RNN layer processes all the tokens on the input sequence iteratively and reutilizes the hidden state of an iteration as the input for generating the hidden state for the next element. The last hidden state of the RNN is passed as input to the first iteration of the decoder. Therefore, when processing a sequence of inputs, the RNN layer generates a sequence of hidden states, with each hidden state being a function of the current input and the previous hidden state. In particular, a bidirectional Gated Recurrent Unit (Bi-GRU) in the RNN layer of the encoder, and a regular GRU in the RNN layer of the decoder are utilized. The structure of the Gated Recurrent Unit (GRU) is presented by [62]. GRU captures the sequential information in the data by maintaining an internal hidden state that evolves over time. It has a gating mechanism that controls the flow of information within the network, making it capable of learning and remembering longer sequences compared to traditional RNNs. The gating mechanism consists of an update gate and a reset gate, which control how much of the previous hidden state should be retained and how much new information should be incorporated in each time step. A Bi-GRU is a variant of the GRU architecture that processes input sequences in both forward and backward directions. The input sequence is fed into two separate GRU layers, one processing the input sequence in the forward direction and the other processing the input sequence in the backward direction. The outputs of the two layers at each time step are concatenated and passed to the next layer or used as the final output.
@@ VI-B Dataset
Figure 6: Spatial distribution of GPS error aggregated per segment.
@@ VI-D Implementation Detail
After evaluating different sets, model hyperparameters are configured as follows. The embedding dimension ($d_{emb}$) is set to 512, while the MLP intermediate dimension ($d_{ff}$) is set to 2048. We use 8 transformer blocks for both the encoder and decoder, and each multi-head attention block employs 16 linear transformations, resulting in $d_{k}=32$. During training, we utilize the Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function. The batch size is fixed at 200, and the model undergoes training for a total of 32 epochs. For the RNN model, the embedding dimension is set to 1024. We employ the Adam optimizer with a learning rate of 0.0005 for both the encoder and decoder. Similarly, the batch size is fixed at 200, and the model undergoes training for a total of 32 epochs reaching to the point that the loss function converges. We implemented the model in Python using the PyTorch library and trained it on a machine equipped with NVIDIA RTX 6000 GPUs.
@@ VIII Conclusion
In conclusion, this paper has introduced and demonstrated the effectiveness of a deep-learning model, specifically the transformer-based encoder-decoder architecture, as a surrogate for advanced map-matching algorithms. This model has shown remarkable capabilities in addressing the challenges posed by noisy GPS data in urban road networks. The encoder in the encoder-decoder framework, is capabale in encoding the sequence of noisy GPS points and capturing their autoregressive behavior and spatial correlations. By associating these representations with road network features, the decoder efficiently transforms them into a sequence of road segments, effectively mapping noisy GPS data to navigated routes. Through rigorous training and evaluation on GPS traces collected in Manhattan, New York, the model has achieved the highest accuracy rate of 80% compared to the accuracy of the RNN-based encoder-decoder models trained in the urban networks of China reported in the literature. This level of performance aligns with the success of transformer-based encoder-decoder models widely used in natural language processing, further underscoring the model’s promise in translating complex and noisy GPS data into accurate navigational routes within urban road networks. In conclusion, this study has fulfilled its primary objective by developing a deep learning surrogate model that can replicate the complex and computationally intensive processes of advanced map-matching algorithms. This surrogate model offers an efficient alternative to traditional matching methods, significantly enhancing both the speed and accuracy of map-matching tasks. Moreover, its ability to autonomously capture underlying mobility patterns and noise structures within trajectory data contributes to more precise and robust matching results. Ultimately, this research marks a substantial advancement in map-matching techniques, providing a powerful and efficient solution for the identification and tracking of trajectories within intricate and densely populated road networks. The transformer-based encoder-decoder model’s success in this context has promising implications for various applications involving location-based services and urban mobility analysis.

### v4

@@ V-B3 Transformer-based surrogate map-matching
The configuration of the porposed transformer-based model is set as follows. The embedding dimension ($d_{emb}$) is set to 512, and the MLP intermediate dimension ($d_{ff}$) is set to 1024. We use $N=4$ transformer blocks for both the encoder and decoder, and each multi-head attention block employs $h=4$ linear transformations, resulting in $d_{k}=128$. The Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function is utilized. The batch size is set to 32, and the model trained for a total of 200 epochs. The inference time for map-matching using the constructed transformer model varies with trajectory length, with an average inference time of 0.30s and a standard deviation of 0.16s for trajectories ranging from 6 to 204 road segments.
@@ VI-A2 RNN-based encoder-decoder with attention mechanism
The main components of the RNN-based encoder-decoder architecture are two layers: the embedding layer and the RNN layer. Therefore, the main difference between the RNN-based encoder-decoder with the transformer-based encoder-decoder is its recurrent layer versus the self-attention layer in the transformer. The architecture of the RNN-based encoder-decoder model can be found in [68]. In an RNN-based encoder-decoder, during the encoding, the RNN layer processes all the tokens on the input sequence iteratively and reutilizes the hidden state of an iteration as the input for generating the hidden state for the next element. The last hidden state of the RNN is passed as input to the first iteration of the decoder. Therefore, when processing a sequence of inputs, the RNN layer generates a sequence of hidden states, with each hidden state being a function of the current input and the previous hidden state. In particular, a bidirectional Gated Recurrent Unit (Bi-GRU) in the RNN layer of the encoder, and a regular GRU in the RNN layer of the decoder are utilized. The structure of the Gated Recurrent Unit (GRU) is presented by [69].
GRU captures the sequential information in the data by maintaining an internal hidden state that evolves over time. It has a gating mechanism that controls the flow of information within the network, making it capable of learning and remembering longer sequences compared to traditional RNNs. The gating mechanism consists of an update gate and a reset gate, which control how much of the previous hidden state should be retained and how much new information should be incorporated in each time step. A Bi-GRU is a variant of the GRU architecture that processes input sequences in both forward and backward directions. The input sequence is fed into two separate GRU layers, one processing the input sequence in the forward direction and the other processing the input sequence in the backward direction. The outputs of the two layers at each time step are concatenated and passed to the next layer or used as the final output.
@@ VI-B Dataset
Figure 9: Spatial distribution of GPS error (in meters) aggregated per road segment.
@@ VI-D Implementation Detail
After evaluating different sets, model hyperparameters are configured as follows. The embedding dimension ($d_{emb}$) is set to 512, and the MLP intermediate dimension ($d_{ff}$) is set to 2048. We use 8 transformer blocks for both the encoder and decoder, and each multi-head attention block employs 16 linear transformations, resulting in $d_{k}=32$. During training, we utilize the Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function. The batch size is fixed at 200, and the model undergoes training for a total of 32 epochs. For the RNN model, the embedding dimension is set to 1024. We employ the Adam optimizer with a learning rate of 0.0005 for both the encoder and decoder. Similarly, the batch size is fixed at 200, and the model undergoes training for a total of 32 epochs reaching to the point that the loss function converges. We implemented the model in Python using the PyTorch library and trained it on a machine equipped with NVIDIA RTX 6000 GPUs.
@@ VIII Conclusion
This study introduces a context-aware deep-learning-based trajectory map-matching framework that predicts the most likely navigated path as an ordered sequence of road segments from noisy and sparse GPS data. Traditional methods rely on geometric proximity, topological features, and simplified assumptions, such as uniform noise distribution and shortest-path heuristics for route estimation. These approaches often use probabilistic models or handcrafted rules, which may not generalize well across diverse urban environments. Unlike conventional methods, large-scale neural networks can infer contextual factors such as non-uniform noise patterns, driver trajectory behavior, and complex road network structures from real-world data. By leveraging large-scale trajectory datasets, these models capture spatial uncertainties and mobility patterns, improving the accuracy and robustness of trajectory map-matching.
Building on the success of large language models, large-scale neural network-based foundation models are now accelerating various scientific domains. Examples include Aurora, trained on large-scale weather and climate data [71], and ChemBERTa, designed for molecular data to aid drug discovery [25]. Recent studies have also envisioned a shift toward context-aware trajectory-powered foundation models [24, 72] for accelerating downstream mobility tasks.
In line with this vision, this study presents an NLP-enabled map-matching framework by reframing map-matching as a sequence-to-sequence translation task, similar to those in NLP. The model adopts a transformer-based encoder-decoder architecture, widely used in natural language processing. By tokenizing noisy GPS trajectories and leveraging self-attention mechanisms, the model learns context-aware representations of driver trajectory behavior, road network structures, and environmental factors that contribute to noise variations. This approach enables the model to effectively translate raw GPS points into their most likely navigated routes, capturing spatial dependencies and contextual information beyond traditional heuristic-based methods. This approach aligns with the trajectory-as-language perspective presented by [26], which seeks to replace traditional, rule-based trajectory analysis with learning-based sequence models that generalize across different urban environments. Similar to how large language models capture latent linguistic structures without predefined rules, this method allows the model to learn spatial correlations, mobility preferences, and GPS noise distributions directly from trajectory data.
The effectiveness of the developed framework is demonstrated through evaluation on real-world GPS traces in Manhattan, New York. The model improved trajectory map-matching accuracy by leveraging both local road network features and global contextual dependencies, serving as a practical step toward trajectory-powered foundation models.
For deep learning-based map-matching frameworks successful out-of-domain generalization requires accounting for variability in road network, built environment, and local driving behaviors. Pretraining on multi-city trajectory datasets and transfer learning [73] can further enhance the model’s ability to generalize to trajectory analysis in different city’s road networks while preserving core spatial and temporal dependencies that drive human mobility. Expanding the training data to include cities with diverse network structures, mobility patterns, and GPS noise distributions could transform this model into a scalable, region-agnostic trajectory analysis tool, contributing to the broader vision of a universal trajectory foundation model for urban mobility [24].
Scaling this model to large cities and multi-city applications requires addressing the increasing complexity of road networks, particularly the growth in unique grid locations ($G$) and road segments ($R$). This scaling challenge is analogous to vocabulary expansion in NLP models, where handling larger vocabularies necessitates techniques such as adaptive embeddings [74, 75], Mixture of Softmax Experts (MoE) [76], language-clustered vocabularies [77], and hierarchical location encoding [78, 79]. Further investigation is needed to understand the relationship between road network size and model performance, particularly in exploring whether similar NLP strategies can enhance trajectory analysis in large-scale road networks. While NLP models provide a strong conceptual foundation, their direct adaptation to trajectory modeling may introduce unique challenges. A deeper exploration within the context of trajectory data is essential for developing scalable, multi-city trajectory models capable of generalizing across diverse urban landscapes.
Future research in NLP-enabled geospatial trajectory analysis is needed to explore the feasibility and effectiveness of these strategies in managing large-scale road networks. While multilingual NLP models provide a valuable conceptual foundation, their direct adaptation to trajectory modeling introduces unique challenges, such as spatial dependencies, network topology constraints, and variations in real-world mobility patterns. Investigating region-aware tokenization, hierarchical location encoding, and adaptive attention mechanisms in the context of trajectory data will be critical to developing scalable, multi-city trajectory models capable of generalizing across diverse urban landscapes.